# GAMA: Generative Agents for Multi-Agent Autoformalization

### v2

@@ Abstract
Multi-agent simulations facilitate the exploration of interactions among both natural and artificial agents. However, modelling real-world scenarios and developing simulations often requires substantial expertise and effort. To streamline this process, we present a framework that enables the autoformalization of interaction scenarios using agents augmented by large language models (LLMs) utilising game-theoretic formalisms. The agents translate natural language descriptions of interactions into executable logic programs that define the rules of each game, ensuring syntactic correctness through validation by a solver. A tournament simulation then tests the functionality of the generated game rules and strategies. After the tournament, if a ground truth payoff matrix is available, an exact semantic validation is performed. We evaluate our approach on a diverse set of 110 natural language descriptions exemplifying five $2\times 2$ simultaneous-move games, achieving 100% syntactic and 76.5% semantic correctness in the generated game rules for Claude 3.5 Sonnet, and 99.82% syntactic and 77% semantic correctness for GPT-4o. Additionally, we demonstrate high semantic correctness in autoformalizing gameplay strategies. Overall, the results highlight the potential of autoformalization to leverage LLMs in generating formal reasoning modules for decision-making agents.
@@ 1 Introduction
Multi-agent simulations provide insights into diverse phenomena, ranging from individual human interactions and organizational dynamics to biological processes, robotics, and interactions among artificial agents. However, modelling real-world scenarios and implementing corresponding simulations typically require substantial human expertise and effort. Consider, for instance, the challenge of coordinating autonomous vehicles: creating a simulation model involves specifying the rules of interaction, the goals of each agent, the information available to them, and the strategies they may employ, among other factors. Once developed, this model must then be implemented. The emergence of Large Language Models (LLMs) enables a novel method for streamlining these tasks through autoformalization. We demonstrate it using scenarios that can be modelled as $2\times 2$ simultaneous-move games.
Although LLMs have been evaluated as agents in game-theoretic contexts Akata et al. (2023); Fan et al. (2023); Guo (2023); Lorè and Heydari (2023), their inherent limitations—such as hallucinations Achiam et al. (2023) and logical or arithmetical errors Imani et al. (2023)—constrain their effectiveness as rational decision—makers in simulations. Rather than relying on LLMs as autonomous agents, a more promising approach is to harness their strengths in format translation, particularly in converting natural language into formal representations. This process, known as autoformalization Wu et al. (2022), has shown promise in formalizing mathematics Wu et al. (2022); He-Yueya et al. (2023); Jiang et al. (2022) and logic Yang et al. (2023); Pan et al. (2023); Cosler et al. (2023); Chen et al. (2023); Feng et al. (2023). The resulting formal representations can be utilized in combination with subsymbolic models within a neuro-symbolic framework. This approach, as demonstrated by recent advances such as AlphaGeometry, has the potential to achieve human-like problem-solving abilities in challenging domains like geometry Trinh et al. (2024).
In this work, we introduce GAMA: Generative Agents for Multi-Agent Autoformalization, a framework that leverages autoformalization to generate formal representations of strategic interactions. These representations serve as the reasoning module of agents deployed in simulations. This research builds upon our prior work Mensfelt et al. (2024), where we presented a basic framework for autoformalizing natural language interaction scenarios. In this paper, we integrate the previously developed autoformalization module into a modular agent architecture within a simulation framework. The main contributions of this work are as follows:
@@
    An agent model with formal game and strategy representation: We propose an agent model that incorporates formal representations of both games and strategies, along with the autoformalization module from our earlier work. Game rules and strategies can either be autoformalized or provided using predefined rules.
@@
    Three-level agent validation: We introduce a validation mechanism for agents at three levels: (1) syntactic correctness, using formal solvers; (2) functionality, by testing the code in the simulation; and (3) semantic correctness, by comparing target and achieved payoffs (when a ground truth payoff matrix is available) or verifying compliance with game type constraints (when a ground truth game type is provided).
@@
    A simulation framework for autoformalized scenarios: We present a simulation framework for autoformalized interaction scenarios, enabling automated validation and facilitating the exploration and comparison of different strategies for specific scenarios.
@@
    Experimental evaluation: We evaluate the framework, demonstrating high levels of syntactic and semantic correctness for scenarios modelled as $2\times 2$ simultaneous-move games. Additionally, we assess the agents’ ability to autoformalize five distinct strategies.
@@ 2.1 Game Theory
Game theory is an analytical framework for strategic interactions, where games are composed of four main components Rasmusen (2006):
  • Players: players consistently and willingly act towards maximising their utility, in full awareness of the situation they are involved in;
  • Actions: mappings between states and outcomes, carried out players Osborne and Rubinstein (1994);
  • Payoffs: payoffs are quantitative measures over the consequent outcomes of action tuples chosen by players Gibbons (1992);
  • Information: game information describes a player’s knowledge about the available actions, historical moves, and the player’s image of its opponents Osborne and Rubinstein (1994).
@@
In the case of symmetric games, the payoffs solely depend on the chosen strategies, while with asymmetric games, the outcomes yield different payoffs based on the player’s identity. In our experiments, the Prisoner’s Dilemma, Hawk-Dove, and Stag-Hunt are symmetric, and we classify actions U and L as “Cooperate” denoted by $C$ and actions D and R as “Defect” denoted by $D$. This allows us to use the terminology from Axelrod’s tournaments Axelrod and Hamilton (1981) to define these games in terms of the four outcomes: $\mathbf{T}:(D,C),\mathbf{R}:(C,C),\mathbf{P}:(D,D),\mathbf{S}:(C,D)$.
  • Prisoner’s Dilemma is characterised by a dominant temptation to defect, represented by the payoff relation $T>R>P>S$,
  • Hawk-Dove entails switched relative payoffs between the punishment for mutual defection and the sucker’s payoff – $T>R>S>P$,
  • Stag Hunt involves a relatively high reward for mutual cooperation over the temptation to defect – $R>T>P>S$.
@@
  • Battle of the Sexes is a coordination game, where both players prefer an agreement but have different preferences over the coordinated outcomes:
    • row: $W>Z>\{X,Y\}$,
@@
  • Matching Pennies: coordinated outcomes reward the row player, while mismatched actions benefit the column player:
    • row: $\{W,Z\}>\{X,Y\}$,
@@ 2.2 Large Language Models
The rapid advancement of natural language processing (NLP), driven by transformer architectures Gillioz et al. (2020) and pre-trained models Qiu et al. (2020), has led to the emergence of Large Language Models (LLMs). State-of-the-art (SOTA) LLMs continue to evolve rapidly, primarily due to the increasing scale of model parameters and training datasets Zhao et al. (2023).
In this work, we employ OpenAI’s GPT-4 Omni OpenAI (2024), a multimodal SOTA model capable of processing text, images, and audio. It has demonstrated strong performance on reasoning benchmarks Wang and Zhao (2024). Likewise, we utilize Anthropic’s Claude 3.5 Sonnet, another high-performing SOTA model, which excels across multiple benchmarks Anthropic (2024).
@@ 2.3 General Game Playing
General game playing Genesereth et al. (2005) focuses on developing intelligent systems that explicitly represent the rules of arbitrary new games and learn to play them autonomously, without human intervention. The Game Description Language (GDL) has been proposed as a formal, machine-processable language for describing the rules of arbitrary games Love et al. (2006). GDL focused on information games only, so it was extended in GDL-II Thielscher (2010) to cover n-player games with incomplete information and games in extensive normal form Thielscher (2011).
A challenge with GDL systems is that learning without human guidance demands complex reasoning. Players must infer the possible actions of others, effectively analysing hypothetical game scenarios before making decisions. Action formalisms like the classical Situation Calculus McCarthy and Hayes (1981) have been developed for precisely this purpose. Formal schemes and inference methods are readily available for Situation Calculus Giacomo et al. (2010); Lesperance et al. (2024), while their deployment in general game playing presupposes a translation from GDL into existing, suitably expressive action languages. One such scheme Schiffel and Thielscher (2011) shows how to fully embed GDL-II into a version of the Situation Calculus based on knowledge fluents Scherl and Levesque (2003). Our game solver is inspired by these previous works, to support light GDL specifications that act as target representation for the autoformalization.
@@ 3.1 Overview
Figure 1: GAMA framework: each agent’s game rules and strategy can be predefined, autoformalized, or loaded from an existing agent. The initial agent pool undergoes syntactic validation, and agents that fail to produce syntactically valid code within a defined number of attempts are eliminated (denoted by a crossed circle). The remaining agents are validated in the tournament and subsequently undergo semantic validation.
The GAMA framework introduced in this work enables the creation of modular autoformalizing agents and supports tournament play using these agents (Fig. 1). An agent consists of an autoformalization module and a logic programming component implemented in Prolog, which is divided into two modules: a Game module, containing both game-independent and game-dependent rules, and a Strategy module, which describes the actions to select against an opponent. Together, these modules form a solver that encapsulates the game logic and determines the agent’s next move. A Python wrapper facilitates the agent’s interaction with the external environment, manages the game history (tracking each agent’s moves, opponent moves, and payoffs), and updates the solver’s state. This modular design allows for easy substitution of game-dependent rules and strategies, either through predefined configurations or by autoformalizing natural language descriptions. The source code and log files generated during the evaluation are available at [^1] https://github.com/dicelab-rhul/GAMA.
3.2 Solver
The solver comprises a game-independent part that represents the rules for any extensive-form game, a game-dependent part that defines the rules of a particular game using the predicates from the game-independent part, and a strategy part to select an action. Our solver is implemented in Prolog Körner et al. (2022); Wielemaker et al. (2012)—a declarative programming language well-suited for high-level problem definitions—but it can easily be substituted with another formal solver. The game-independent and game-dependent parts were described in detail in Mensfelt et al. (2024). To enable the agent incorporating the solver to play, we extended it with the strategy part. Additionally, we implemented semantic validation in Prolog using validity constraints. The strategy part and constraint-checking mechanism are described below.
3.2.1 Strategy
select(P, O, S, M) if
not holds(last_move(O, _LMo), S) and
holds(default_move(P, M), S).
select(P, O, S, Mo) if
holds(last_move(O, Mo), S).
Here, a player P selects the default move first (typically to cooperate), otherwise it mirrors the opponent’s move in the last round. holds/2 is based on the implementation of the Situation Calculus McCarthy and Hayes (1981). The strategy relies on specific predicates – default_move/2 and last_move/2 – which are required from the game-dependent part of the solver. More sophisticated strategies can be expressed.
3.2.2 Validity checking
We check that the auto-formalized payoff matrix satisfies the structural and content constraints of the specified game (as outlined in Sect. 2.1), by defining validity constraints. The listing below specifies the payoff matrix validity constraints for the Prisoner’s Dilemma.
valid_pd_payoffs(T,R,P,S,C,D) if
payoff(C,C,R,R) and
payoff(C,D,S,T) and
T>R and
payoff(D,C,T,S) and
payoff(D,D,P,P) and
R>P and
P>S.
3.3 Autoformalization Module
@@
To perform autoformalization, a large language model is employed. The framework supports the use of any language model by providing an abstract interface. Autoformalization of a natural language description of a game or strategy is achieved through one-shot prompting, using game-dependent predicates for the Prisoner’s Dilemma or the tit-for-tat strategy as examples, respectively. After generating the code, its syntactic correctness is validated. If errors are detected, the LLM receives feedback, including the erroneous lines, error messages, and instructions for correcting the predicates. The max_attempts parameter specifies the maximum number of correction attempts. The overview of autoformalization algorithm is presented in Listing 1 Mensfelt et al. (2024).
3.4 Simulation
The framework is employed in this work in two primary modes: validation and simulation. These modes facilitate the evaluation of both the semantic correctness of generated programs and the effectiveness of strategies within various game scenarios.
In validation mode, the framework ensures the semantic correctness of generated programs by comparing their performance against predefined criteria. Once a scenario is autoformalized, an agent is created and provided with a specified strategy. The agent then plays against a clone of itself for a set number of rounds. If a ground truth payoff matrix is available, the framework performs semantic validation by verifying that the payoff for each unique combination of moves matches the target payoff for that combination. Additionally, if the ground truth type of the game is known, the framework enforces constraint checking to confirm that the payoff matrix adheres to all specified constraints for the given game.
In simulation mode, the framework evaluates and compares strategies to determine their effectiveness within a game scenario. Agents equipped with different strategies compete in a round-robin tournament, where each agent plays against every other. By simulating multiple rounds, the framework identifies the strategy that performs best on average, offering insights into the relative strengths and weaknesses of competing approaches.
The parameters of the tournament are as follows:
  • Agent Pool: Contains all agents participating in the tournament. Each agent can be equipped with game rules and strategies either through autoformalization or predefined configurations.
  • Number of Rounds: Specifies how many rounds each agent will play.
  • Match Maker: A function that determines the pairing of agents for matches.
  • Target Payoffs: Optional parameter used for semantic validation.
The framework provides two methods for determining the tournament’s outcome:
@@ 3.4 Simulation
    Selecting agents whose total payoff matches their target payoff, applicable when target payoffs are specified (validation).
@@
    Selecting agents that achieve the highest total payoff, suitable for assessing the success of strategies (simulation).
Furthermore, the customizable Match Maker function allows for experiments with tailored matching protocols

### v3

# Generative Agents for Multi-Agent Autoformalization of Interaction Scenarios

@@ Abstract
Multi-agent simulations are versatile tools for exploring interactions among natural and artificial agents, but their development typically demands domain expertise and manual effort. This work introduces the Generative Agents for Multi-Agent Autoformalization (GAMA) framework, which automates the formalization of interaction scenarios in simulations using agents augmented with large language models (LLMs). To demonstrate the application of GAMA, we use natural language descriptions of game-theoretic scenarios representing social interactions, and we autoformalize them into executable logic programs defining game rules, with syntactic correctness enforced through a solver-based validation. To ensure runtime validity, an iterative, tournament-based procedure tests the generated rules and strategies, followed by exact semantic validation when ground truth outcomes are available. In experiments with 110 natural language descriptions across five $2\times 2$ simultaneous-move games, GAMA achieves 100% syntactic and 76.5% semantic correctness with Claude 3.5 Sonnet, and 99.82% syntactic and 77% semantic correctness with GPT-4o. The framework also shows high semantic accuracy in autoformalizing agents’ strategies.
@@ 1 Introduction
Multi-agent simulations (MAS) provide insights into diverse phenomena, ranging from individual human interactions and organizational dynamics to biological processes, robotics, and interactions among artificial agents. However, modelling real-world scenarios and implementing corresponding simulations typically require substantial human expertise and effort. Recent advances in large language models (LLMs) and autoformalization techniques [37] offer a promising new approach to streamlining these tasks. In this work, we demonstrate this approach using interaction scenarios modelled as $2\times 2$ simultaneous-move games.
LLMs have been increasingly explored as agents in interaction scenarios and game-theoretic contexts, though often under a weak notion of agency [2, 7, 14, 19]. Furthermore, their inherent limitations—such as hallucinations [1] and logical or arithmetic errors [16]—undermine their reliability as rational decision-makers in simulations. Addressing this challenge, autoformalization shifts the role of LLMs from decision-making to format translation, converting natural language interaction descriptions into formal representations. This leverages one of LLMs’ core strengths, with autoformalization showing considerable promise in formalizing mathematics [37, 15, 17] and logic [38, 26, 6, 5, 8].
In multi-agent systems, such formalized interaction descriptions can aid and, often act, as decision-making modules, reducing reliance on the LLM’s reasoning correctness during runtime. Moreover, this formal representation can be iteratively refined using automatic reflection and refinement mechanisms [21, 29, 32, 13], enabling continuous improvement via feedback loops. This shift from natural language reasoning to formal, verifiable representations establishes a more robust foundation for automating LLM-driven agent design.
Our prior work [23], introduced an autoformalization module capable of converting interaction descriptions into syntactically validated code. The Generative Agents for Multi-Agent Autoformalization (GAMA) incorporates the previous work as a component of the simulation framework, while additionally providing runtime and semantic validation and enabling the generation of formal representations of both strategic interactions and game strategies. These formalizations act as reasoning modules for agents deployed in simulations, enabling comparison of different strategies for given interactions.
The main contributions of this work are as follows:
@@
    Agent model with formal game and strategy representations: We propose an agent architecture that integrates formalized representations of games and strategies alongside the autoformalization module. Game rules and strategies can be autoformalized or supplied as predefined specifications.
@@
    Three-level agent validation: We introduce a multi-level validation pipeline for agents, including (1) syntactic validation using formal solvers, (2) runtime validation through simulation-based testing, and (3) semantic validation based on ground truth outcomes (when available) or compliance with game-specific constraints.
@@
    Simulation framework for autoformalized scenarios: We present a simulation environment that supports automated validation and facilitates the exploration and comparison of strategies formalized from natural language descriptions.
@@
    Experimental evaluation: We empirically evaluate GAMA, demonstrating high syntactic and semantic correctness across 110 scenarios modelled as $2\times 2$ simultaneous-move games. Furthermore, we assess the agents’ ability to autoformalize and execute five distinct gameplay strategies.
The remainder of this paper is organized as follows. Section 2 introduces the background and preliminaries. Section 3 presents the GAMA framework. Section 4 details the experimental setup and evaluation methodology. Section 5 reports the experimental results. Finally, Section 6 concludes the paper and outlines directions for future work.
@@ 2.1 Game Theory
Game theory is an analytical framework for strategic interactions, where games are composed of four main components [28]:
Players
players consistently and willingly act towards maximising their utility, in full awareness of the situation they are involved in.
Actions
mappings between states and outcomes, carried out by players [25].
Payoffs
payoffs are quantitative measures over the consequent outcomes of action tuples chosen by players [11].
Information
game information describes a player’s knowledge about the available actions, historical moves, and the player’s image of its opponents [25].
@@
In the case of symmetric games, the payoffs solely depend on the chosen strategies, while with asymmetric games, the outcomes yield different payoffs based on the player’s identity. In our experiments, the Prisoner’s Dilemma, Hawk-Dove, and Stag-Hunt are symmetric, and we classify actions U and L as “Cooperate” denoted by $C$ and actions D and R as “Defect” denoted by $D$. This allows us to use the terminology from Axelrod’s tournaments [4] to define these games in terms of the four outcomes: $\mathbf{T}:(D,C),\mathbf{R}:(C,C),\mathbf{P}:(D,D),\mathbf{S}:(C,D)$.
Prisoner’s Dilemma
is characterised by a dominant temptation to defect, represented by the payoff relation $T>R>P>S$.
Hawk-Dove
entails switched relative payoffs between the punishment for mutual defection and the sucker’s payoff: $T>R>S>P$.
Stag Hunt
involves a relatively high reward for mutual cooperation over the temptation to defect: $R>T>P>S$.
@@
Battle of the Sexes
is a coordination game, where both players prefer an agreement but have different preferences over the coordinated outcomes:
  • Row: $W>Z>\{X,Y\}$.
@@
Matching Pennies
coordinated outcomes reward the row player, while mismatched actions benefit the column player:
  • Row: $\{W,Z\}>\{X,Y\}$.
@@ 2.2 Large Language Models
The rapid advancement of natural language processing (NLP), driven by transformer architectures [12] and pre-trained models [27], has led to the emergence of Large Language Models. State-of-the-art (SOTA) LLMs continue to evolve rapidly, primarily due to the increasing scale of model parameters and training datasets [39].
In this work, we employ OpenAI’s GPT-4 Omni [24], a multimodal SOTA model capable of processing text, images, and audio. It has demonstrated strong performance on reasoning benchmarks [35]. Likewise, we utilize Anthropic’s Claude 3.5 Sonnet, another high-performing SOTA model, which excels across multiple benchmarks [3].
@@ 2.3 General Game Playing
General game playing [9] focuses on developing intelligent systems that explicitly represent the rules of arbitrary new games and learn to play them autonomously, without human intervention. The Game Description Language (GDL) has been proposed as a formal, machine-processable language for describing the rules of arbitrary games [20]. GDL focused on information games only, so it was extended in GDL-II [33] to cover n-player games with incomplete information and games in extensive normal form [34].
A challenge with GDL systems is that learning without human guidance demands complex reasoning. Players must infer the possible actions of others, effectively analysing hypothetical game scenarios before making decisions. Action formalisms like the classical Situation Calculus [22] have been developed for precisely this purpose. Formal schemes and inference methods are readily available for Situation Calculus [10, 18], while their deployment in general game playing presupposes a translation from GDL into existing, suitably expressive action languages. One such scheme [31] shows how to fully embed GDL-II into a version of the Situation Calculus based on knowledge fluents [30]. Our game solver is inspired by these previous works, to support light GDL specifications that act as target representation for the autoformalization.
@@ 3.1 Overview
Figure 1: Overview of the GAMA framework. Dashed lines indicate optional control flows. Games and strategies can be specified either through predefined rules or natural language descriptions, or optionally by loading a predefined agent. Semantic validation is performed when ground truth outcomes are available. The Validator may employ a formal solver.
Fig.1 presents an overview of the GAMA framework. The framework is implemented in Python and Prolog [36], though its modular architecture supports substituting alternative solvers. The source code and log files generated during the evaluation are available at [^1] https://github.com/dicelab-rhul/GAMA.
Each agent comprises an autoformalization module and a logic programming component, which is further divided into two modules: a Game module, containing both game-independent and game-specific rules, and a Strategy module, which specifies the agent’s action selection logic against an opponent. Together, these modules form a solver that encapsulates the game logic and determines the agent’s next move. A Python wrapper enables interaction with the external environment, manages game history (including each agent’s moves, opponent actions, and payoffs), and updates the solver’s internal state accordingly.
3.2 Agents’ Initialization and Syntactic Validation
@@
To perform autoformalization, an agent employs a large language model (LLM) through a unified abstract interface, allowing compatibility with any LLM. Natural language descriptions of games or strategies are formalized using one-shot prompting, with examples of game-dependent predicates (e.g., for the Prisoner’s Dilemma or tit-for-tat strategy) provided in the prompt. Once the code is generated, its syntactic correctness is validated. If errors are detected, the LLM receives feedback containing the erroneous lines, error messages, and instructions for correcting the predicates. The process is repeated up to a limit defined by the max_attempts parameter. An overview of the autoformalization module [23] is shown in Listing 1.
3.3 Simulation and Runtime Validation
After generating a pool of syntactically correct agents, the framework operates in two primary modes: validation and strategy evaluation. These modes enable assessment of both the runtime correctness of generated programs and the effectiveness of strategies across various game scenarios.
The key parameters for tournament-based simulation are as follows:
Agent Pool
All participating agents, initialized with game rules and strategies obtained either through autoformalization or predefined configurations.
Number of Rounds
The number of rounds each agent pair will play.
Match Maker
A customizable function that defines how agents are paired for matches.
Target Payoffs
(Optional) Used for semantic validation when ground truth payoffs are available.
Tournament outcomes can be determined in two ways:
@@ 3.3 Simulation and Runtime Validation
    Selecting agents that achieve the highest total payoff, reflecting strategy effectiveness.
@@
    Selecting agents whose total payoff matches the target payoff (if available), used primarily for validation.
In strategy evaluation mode, the framework compares strategies based on their performance within game scenarios. Agents equipped with different strategies typically compete in a round-robin tournament, where each agent plays against every other agent for a specified number of rounds. However, alternative competition formats can be defined using a custom Match Maker function. By simulating multiple matches, the framework identifies which strategies perform best on average, providing insights into their relative strengths and weaknesses.
3.4 Semantic Validation
Semantic validation is performed when ground truth outcomes are available. Here, “semantic correctness” refers to the correspondence between the generated code and the input natural language description. For example, formalized code for the Prisoner’s Dilemma that is syntactically valid and executes correctly in simulation, but encodes an incorrect payoff matrix, would be considered semantically incorrect.
In semantic validation mode, when specific target outcomes are provided, the framework performs exact validation. Once a scenario is autoformalized and an agent is instantiated with the given strategy, the agent plays against a clone of itself for a fixed number of rounds. The resulting outcomes for all action tuples are then compared directly to the target outcomes to assess semantic correctness.
If specific target outcomes are not available but the game type is known (e.g., it is known that the scenario represents a Prisoner’s Dilemma, but the exact payoff matrix is unknown), the framework performs constraint-based validation. In this case, it verifies that the generated payoff matrix satisfies all necessary properties (such as symmetry or payoff dominance relations) defined for that game type (Section 3.5.4). This ensures that the generated logic faithfully captures the intended structure of the game.
3.5 Solver
Our solver is based on logic programming, as described in [23], but repeated here for completeness. It comprises of a game-independent part representing the rules for any normal-form game, a game-dependent part defining the rules of a particular game using the predicates from the game-independent part, and a set of auxiliary predicates completing the definition of that particular game. We follow standard Prolog conventions to interpret a game: variables are indicated by uppercase letters, while predicates and function symbols by lowercase. An underscore ’_’ represents a variable whose value is ignored within a definition. The game state is then a situation, initially denoted by a constant (e.g. s0). The binary function do(M, S) denotes the situation resulting from the executing move M in situation S, consistent with the Situation Calculus.
3.5.1 Game-independent part
We define all legal transitions of an extended-form game from an initial situation S to a final situation F as follows:
game(F, F) if
final(F).
game(S, F) if
not final(S) and
legal(M, S) and
game(do(M, S), F).
The game terminates when the final game situation F is reached. Otherwise, in a non-final situation S, the game accepts a legal move M, and the game continues in the next do(M,S) situation, until the final situation F is reached. To determine what holds in each legal situation, we use Situation Calculus, represented as:
holds(F, S) if
initially(F, S).
holds(F, do(M, S)) if
effect(F, M, S).
holds(F, do(M, S)) if
holds(F, S) and
not abnormal(F, M, S).
holds/2 here is similar to true/1 in GDL, but with an additional parameter of the situation in which the fluent is true. It states that a fluent F holds in the initial situation (init/1 in GDL), a new fluent F is initiated by the effects of a move M executed in a situation S (next/1 in GDL), and a fluent F persists after a move is made, provided it is not abnormal; abnormal fluents do not persist (implicit in GDL). We also use rules of the form:
finally(F, S) if Conditions.
to return derived fluents F describing the result of the game, when the Conditions hold in the final situation S.
3.5.2 Game-dependent part
To represent a specific game we need to define game-dependent predicates for the initial state initial/1, the legal moves legal/2, what holds in the initial game situation via initially/2, the effects of a move on a situation via effect/3, what stops persisting in a situation after the execution of a move via abnormal/3, the final situation final/1, and the result of the game via finally/2 definitions. To exemplify these definitions, we show how to describe a PD game to our solver. The initial situation s0 is defined as:
initial(s0).
What holds in this initial situation we specify it as:
initially(player(p1), s0).
initially(player(p2), s0).
initially(role(p1,row), s0).
initially(role(p2,col), s0).
initially(control(p1), s0).
initially(control(p2), s0).
These assertions define first the player names represented by unique identifiers (p1 and p2), their roles (p1 is the row player, while p2 is the column player), and then the fact that initially either of them can the game next (via the control/1 fluent,as in GDL). What holds in the initial situation changes as a result of move being made in the game. We represent moves as terms of the form move(P, M), where P is a player, and M is a move. As in PD it is possible for any player to defect (’D’) or cooperate (’C’), we express this as:
possible(move(P,’D’),S) if holds(player(P),S).
possible(move(P,’C’),S) if holds(player(P),S).
It is then legal for a player to make a possible move if they have the control to execute it in the current situation:
legal(move(P, M), S) if
possible(move(P, M), S) and
holds(control(P), S).
As an effect of a legal move M being made by a player P, we record in the next situation that the player did that move (effect/3 is like next/1 in GDL):
effect(did(P, M), move(P, M), S).
Once a legal move is executed, the player loses control, which we specify in our framework as:
abnormal(control(P), move(P, M), S).
When a player loses control cannot play a move until control is passed back. When each player makes a move from the initial situation, the resulting final situation can be determined as:
final(do(M2, do(M1, S))) if
ground(M2) and
ground(M1) and
initial(S).
Assuming the payoff matrix following the structure presented in Table 1 defined as:
payoff(’D’, D’, 1, 1).
payoff(’C’, D’, 0, 5).
payoff(’D’, C’, 5, 0).
payoff(’C’, C’, 3, 3).
the outcome of the game holds information about the actual moves made by the players, and their payoffs.
finally(outcome(P1, M1, U1, P2, M2, U2), S) if
holds(role(P1, row), S) and
holds(did(P1, M1), S) and
holds(role(P2, col), S) and
holds(did(P2, M2), S) and
payoff(M1, M2, U1, U2).
We can then extract specific outcome information, e.g. the player’s utility, through a goal/2 fluent (as in GDL) e.g.:
finally(goal(P1, U1), S) if
finally(outcome(P1, _, U1, _, _, _), S).
finally(goal(P2, U2), S) if
finally(outcome(_, _, _, P2, _, U2), S).
This finalises the definition of PD, enabling us to use the game description above to reason about the game’s dynamics. For instance, if player p1 wanted to analyse the game and identify the actions required to achieve a utility of 5, then it asks the query:
?- game(s0,F), finally(goal(p1,5),F).
Applying the PD payoff matrix, our solver produces two solutions for F, separated by semicolons (;). The result false indicates "no further answers":
F = do(move(p2,’C’),do(move(p1,’D’),s0));
F = do(move(p1,’D’),do(move(p2,’C’),s0));
false.
In the first solution, p1 acts first and p2 second, while in the second solution, the order is reversed. This outcome is expected, as both players have initial control in our game definition, resulting in both action order combinations being shown.
3.5.3 Strategy
select(P, O, S, M) if
not holds(last_move(O, _LMo), S) and
holds(default_move(P, M), S).
select(P, O, S, Mo) if
holds(last_move(O, Mo), S).
Here, a player P selects the default move first (typically to cooperate), otherwise it mirrors the opponent’s move in the last round. holds/2 is based on the implementation of the Situation Calculus [22]. The strategy relies on specific predicates—default_move/2 and last_move/2—which are required from the game-dependent part of the solver. More sophisticated strategies can be expressed.
3.5.4 Constraints validation
We check that the autoformalized payoff matrix satisfies the structural and content constraints of the specified game (as outlined in Sect. 2.1), by defining validity constraints. The listing below specifies the payoff matrix validity constraints for the Prisoner’s Dilemma.
valid_pd_payoffs(T,R,P,S,C,D) if
payoff(C,C,R,R) and
payoff(C,D,S,T) and
T>R and
payoff(D,C,T,S) and
payoff(D,D,P,P) and
R>P and
P>S.